# Unsupervised motion segmentation in one go: Smooth long-term model over a video

### v2

@@ 5.2 Quantitative and comparative evaluation
As shown in Table 2, our LT-MS method provides very convincing results, both for LT-MS-K2 and LT-MS-K4, in the category of unsupervised methods based on optical flow only. Besides, we report a repeatability study in the Appendix. OCLR and DivA demonstrate better performance on the SegTrackV2 dataset. However, as aforementioned, OCLR is not a fully unsupervised method, while DivA leverages RGB images in its conditional decoder. In addition, DivA, along with MoSeg and CIS methods, takes multi-step flows as input between $t$ and in turn $t+1$, $t+2$, $t-1$, $t-2$, and averages the four corresponding predictions to get the final result. SegTrackV2 includes sequences acquired with a poorly controlled handheld camera, which leads to unstable sequences where the contribution of our method is therefore less likely to be emphasized. Failure cases may occur for very complex motion (ripples on the water), or with a particularly strong temporal evolution (an object approaching the camera from afar quickly and closely).
@@
Overall, temporal consistency is properly handled over long temporal periods by our LT-MS method. Beyond segmentation performance, we want to stress that our method is the only one providing by design a coherent segmentation over the sequence, which is a significant added value for downstream tasks, like tracking or dynamic scene interpretation.
5.3 Multi-segment evaluation
We have also performed the evaluation of multiple motion segmentation for a multi-segment setting. Since multiple-motion segmentation is harder than the binary motion segmentation (moving foreground vs background), accuracy scores are expected to decrease for all methods. In Table 3, we report comparative results on the DAVIS2017-motion dataset and on FBMS59. As for the other methods, we performed segmentation with three masks. To this end, we finetuned the LT-MS-K4 network on the DAVIS2016 training set with now three masks ($K=3$). The resulting performance on DAVIS2017-motion is slightly better than ST-MS [4], and far better than MoSeg. There is still the same remark about OCLR status as an unsupervised method. Regarding FBMS59, we report multimask segmentation results with two different metrics. Our LT-MS method outperforms by a large margin DivA [2] that attempted this multi-segment evaluation on FBMS59.
Table 3: Multi-segment evaluation. Evaluation is performed on the video as a whole following the official DAVIS2017 scheme. Reported scores are the average of the individual video scores. $\mathcal{J}\&\mathcal{F}$ is the mean of the two. For FBMS59, the evaluation metrics are bootstrap IoU (bIoU) [2] where each ground-truth mask is mapped to the most likely predicted segment (performed at the frame level), and linear assignment that is a bilinear mapping between the ground-truth and the predicted segments at the sequence level (as in the official DAVIS2017 evaluation). *See caption of Table 2.
Dataset DAVIS2017-motion FBMS59
Method / Scores $\mathcal{J}\&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow bIoU Linear Assignment
@@ 5.4 Qualitative visual evaluation
Fig.3 contains several visual results to demonstrate how our LT-MS method behaves on different situations. We display three result samples obtained on different videos of the benchmarks. Additional results are provided in the Appendix. We can observe that the motion segments are globally accurate. Since our method involves several masks, we can properly handle articulated motions (people2, bmx), deal with the presence of several moving objects in the scene (people2), separate the rider from the bike or motorbike (bmx, motocross-jump), or accomodate motion parallax. Since our method enforces temporal consistency, it can also deal with errors in optical flow estimation or with objects that momentarily stop moving (bmx). We must keep in mind that our actual target is the optical-flow segmentation (OFS) task, even if we evaluate our method on VOS benchmarks. When VOS benchmarks deal with the segmentation of one primary object moving in the foreground, it may occur discrepancies with OFS, which negatively impacts the evaluation scores. The segmentation of additional moving parts, wrong for VOS ground truth, makes sense from the OFS standpoint.
@@ 6 Conclusion
We have designed an original transformer-based unsupervised method for segmenting multiple motions over a video, without any postprocessing. It performs it in one go, which is a more demanding goal, but provides decisive added value regarding for instance tracking or dynamic scene interpretation. Our LT-MS method leverages the ELBO framework for the loss function, and fully acknowledges the temporal dimension of the motion segmentation problem. Our method delivers a stable and consistent OF segmentation throughout long video sequences. It introduces at training time B-splines spatio-temporal parametric motion models over space-time volumes, and a loss term expressing temporal consistency over successive masks while taking care of occlusions. Our transformer-based network can be applied at test time to input volumes of any time length. It can accomodate different choices on the number $K$ of masks with a simple fine-tuning step. Experimental results on several datasets demonstrate the efficiency and the accuracy of our LT-MS method by providing competitive results on several datasets. In addition, it is very fast at test time.
Appendix A Appendix
A.1 Loss function
Here, we motivate the expression of the temporal consistency term $\mathcal{L}_{c}$ of the loss function, given by Eq.7 in the main text.
First, let us recall the main definitions and notations. We consider a volume of optical flow fields $f\in\mathbb{R}^{2\times T\times W\times H}$, the spatial grid $\Omega\in\mathbb{R}^{W\times H}$ and $T$ temporal steps. We assume that we can decompose the flow as a set of $K$ segments, each one exhibiting a coherent motion. Flow vectors within a given segment $k$ can be represented by a smooth parametric motion model parametrized by $\theta_{k}$. $\theta=\{\theta_{k},k=1\cdots K\}$. Variable $z_{i,t}$ conveys the motion segmentation, $z_{i,t}^{k}=1$ if site $(i,t)$ belongs to segment $k$ ($z_{i,t}^{k}=0$ otherwise). $z$ and $\theta$ are latent variables, and $f$ is the observed data.
A.1.1 Prior
We introduce a prior on $z$:
$\forall i\in I,\forall t\in\{1\cdots T\},z_{i,t}\in\{1\cdots K\}$
$p(z)=\prod_{i}p(z_{i,1})\prod_{t=2}^{T}p(z_{i,t}|z_{i,t-1})$
$p(z_{i,t}=k|z_{i,t-1}=l)=\frac{1}{M}\exp(\alpha\mathbb{1}[k=l]+\beta\mathbb{1}[k\neq l])$
$p(z_{i,1}=k)=\frac{1}{K}.$
We also use the notation $g_{i,t}^{k}=g_{\phi}(f)_{i,t}^{k}=q(z_{i,t}=k|f)$.
Using an uninformative prior on $\theta$, we can compute the KL term as:
$KL[q(z,\theta|f)||p(z,\theta)]$ $=KL[q(z|f)||p(z)]$ (10)
$=\mathbb{E}_{q}[\log q(z|f)]-\mathbb{E}_{q}[\log p(z)]$ (11)
$\mathbb{E}_{q}[\log q(z|f)]$ $=\sum_{i,t,k}g_{i,t}^{k}\log g_{i,t}^{k}$ (12)
$\mathbb{E}_{q}[\log p(z)]$ $=\sum_{i}[\mathbb{E}_{q}[\log p(z_{i,1})]+\sum_{t=2}^{T}\mathbb{E}_{q}[\log p(z_{i,t}|z_{i,t-1})]]$ (13)
$=\sum_{i}\sum_{t=2}^{T}\mathbb{E}_{q}[\log p(z_{i,t}|z_{i,t-1})]-I\log(K)$ (14)
$=\sum_{i}\sum_{t=2}^{T}\sum_{k}^{K}\sum_{l}^{L}g_{i,t}^{k}g_{i,t-1}^{l}(\alpha\mathbb{1}[k=l]+\beta\mathbb{1}[k\neq l])+c,$ (15)
where $c=-I(\log(K)+T\log(M))$ is a constant.
We set $\beta=0$, knowing that $\alpha>0$ :
$\mathbb{E}_{q}[\log p(z)]$ $=\alpha\sum_{i}\sum_{t=2}^{T}\sum_{k}^{K}g_{i,t}^{k}g_{i,t-1}^{k}+c.$ (16)
A.1.2 Approximation
$2*g_{i,t}^{k}g_{i,t-1}^{k}=-(g_{i,t}^{k}-g_{i,t-1}^{k})^{2}+(g_{i,t}^{k})^{2}+(g_{i,t-1}^{k})^{2}$ (17)
Thus, we can write:
$\mathbb{E}_{q}[\log p(z)]$ $=\frac{\alpha}{2}\sum_{i}\sum_{t=2}^{T}\sum_{k}^{K}-(g_{i,t}^{k}-g_{i,t-1}^{k})^{2}+(g_{i,t}^{k})^{2}+(g_{i,t-1}^{k})^{2}+c.$ (18)
$\mathbb{E}_{q}[\log p(z)]$ $=\frac{\alpha}{2}(\sum_{i}\sum_{t=2}^{T}\sum_{k}^{K}-(g_{i,t}^{k}-g_{i,t-1}^{k})^{2}$
$+\sum_{i}\sum_{t=1}^{T}\sum_{k}^{K}(g_{i,t}^{k})^{2}+\sum_{i}\sum_{t=2}^{T-1}\sum_{k}^{K}(g_{i,t}^{k})^{2})$ (19)
Thus, we get:
$KL[q(z|f)||p(z)]$ $=\frac{\alpha}{2}\sum_{i}\sum_{t=2}^{T}\sum_{k}(g_{i,t-1}^{k}-g_{i,t-1}^{k})^{2}+\sum_{i}\sum_{t=1}^{T}\sum_{k}g_{i,t}^{k}*(\log(g_{i,t}^{k})$
$-\frac{\alpha}{2}*g_{i,t}^{k})-\frac{\alpha}{2}\sum_{i}\sum_{t=2}^{T-1}\sum_{k}^{K}(g_{i,t}^{k})^{2})+c.$ (20)
$\forall x\in[0,1];\alpha>0:-\frac{\alpha}{2}\leq x(\log x-\frac{\alpha}{2}x)\leq 0$, therefore:
A.1.3 Combining with likelihood
$Elbo(f)$ $=\mathbb{E}_{q(z,\theta|f)}[\log p(f|z,\theta)]-KL(q(z,\theta|f)||p(z,\theta))-c$
$Elbo(f)$ $\geq\mathbb{E}_{q(z,\theta|f)}[\log p(f|z,\theta)]-\frac{\alpha}{2}\sum_{i}\sum_{t=2}^{T}\sum_{k}(g_{i,t}^{k}-g_{i,t-1}^{k})^{2}-c$
As $\forall x,y\in[0,1]^{2}:|x-y|\geq(x-y)^{2}$:
$Elbo(f)$ $\geq\mathbb{E}_{q(z,\theta|f)}[\log p(f|z,\theta)]-\frac{\alpha}{2}\sum_{i}\sum_{t=2}^{T}\sum_{k}|g_{i,t}^{k}-g_{i,t-1}^{k}|-c$ (21)
$Elbo(f)$ $\geq-\sum_{t=1}^{T}\sum_{i\in\Omega}\sum_{k}g_{\phi}(f)_{i,t}^{k}||f(i,t)-\tilde{f}_{S_{n}(\theta_{k})}(i,t)||_{1}/\xi_{f,t}$ (22)
$-\frac{\alpha}{2}\sum_{i,t=2,k}|g_{i,t}^{k}-g_{i,t-1}^{k}|-c$ (23)
This is the lower bound we are maximising, although, in practice we are not accounting a fraction of the sites to deal with occlusion areas.
A.1.4 Experimental validation
In order to evaluate how similar the objective we trained in the paper is to the expression defined in Eq.7 above, we train a model minimizing the following loss function:
$\mathcal{L}=$ $\sum_{t=1}^{T}\sum_{i\in\Omega}\sum_{k}g_{\phi}(f)_{i,t}^{k}||f(i,t)-\tilde{f}_{S_{n}(\theta_{k})}(i,t)||_{1}/\xi_{f,t}$
$-\alpha\sum_{i}\sum_{t=2}^{T}\sum_{k}^{K}g_{i,t}^{k}g_{i,t-1}^{k}+\beta\sum_{t=1}^{T}\sum_{i\in\Omega}\sum_{k}g_{\phi}(f)_{i,t}^{k}\log g_{\phi}(f)_{i,t}^{k}$ (24)
In practice, we use $\alpha=0.5$ and $\beta=0.01$, and we obtain the results reported in Table 4.
Table 4: Comparison of the two losses
Prior / Dataset DAVIS2016 FBMS59 SegTrackv2
Cut Size 10 120 10 120 10 120
Full Model LT-MS-K4 74.8 72.4 61.0 58.2 61.3 60.4
Prior Eq.7 73.5 71.6 57.6 55.6 59.0 57.4
A.2 Datasets
DAVIS2016 comprises 50 videos (for a total of 3455 frames), depicting diverse moving objects. It is split in a training set of 30 videos and a validation set of 20 videos. Only the primary moving object is annotated in the ground truth. The criteria for evaluation on this dataset are the Jaccard score $\mathcal{J}$ (intersection-over-union), and the contour accuracy score $\mathcal{F}$, the higher, the better for both. FBMS59 includes 59 videos (720 annotated frames), and SegTrackV2 14 videos (1066 annotated frames). Both mostly involve one foreground moving object, but sometimes a couple of moving objects. For FBMS59, we use the 30 sequences of the validation set. Although annotations may comprise multiple objects, the VOS community exploits FBMS59 and SegTrackV2 similarily as DAVIS2016, considering a binary ground-truth, by grouping moving objects in the foreground. We follow this practice. However, we also provide a multi-mask evaluation for FBMS59.
A.3 Addition to the ablation study
Fig.4 plots the performance scores for the three ablations when the length of the input flow sequence varies from 10 to 120. Overall, they exhibit comparable behaviour at a certain distance from the full model.
Figure 4: Influence of the length of the input flow sequence (also referred to as cut size on the horizontal axis of the plot) on the three modified versions of our LT-MS-K4 method with comparison with the full model. Left plot: for the DAVIS2016 dataset. Middle plot: for the FBMS59 dataset. Right plot: for the SegTrackV2 dataset. Overall, they exhibit comparable behaviour at a certain distance from the full model.
Visual results reported in Fig.5 clearly demonstrate that the addition of the temporal consistency loss term $\mathcal{L}_{c}$ allows us to get far more consistent segments over time, whether for the background, or the moving objects. For instance, in the fourth example, the foreground moving car is perfectly segmented throughout the sequence along with its wheels, and (small) moving cars in the background. Fig.9 highlights the contribution of the spline-based motion model on the dog video, and its obvious ability to handle motions that do not vary uniformly, as the erratic movement of the dog.
Figure 5: Four groups of qualitative results regarding the ablation of the temporal-consistency loss term. They respectively correspond to the worm video of SegTrackV2, the dogs01 video of FBMS59, and the kite-surf and car-roundabout videos of DAVIS2016. For each group, the first row contains sample images with the segmentation ground-truth, when available at that frame, overlaid in yellow, the second row displays the input flows, the third and fourth rows show the predicted motion segmentations, respectively without and with the temporal consistency loss term. Clearly, this model component allows us to get far more consistent segments over time.
A.4 Optimization implementation
We use Adam optimizer with the following strategy on the learning rate $\alpha$ to train the network (3D Unet and transformer), inspired from the warmup-decay strategy in [5]. We linearly increase it from 00 to $41e-$ 4 for 20 epochs, then, we divide it by two every 40 epochs. The estimation of the motion model parameters through the B-spline approximation at training time is achieved with the Pytorch implementation of L-BGFS [3]. Indeed, as written in eq.2 of the paper, the parametric models induced over time by the splines are a linear combination of quadratic parametric models. Thus, we can optimize eq.6 of the paper to obtain the argmin in eq.9, which is a convex problem.
A.5 Segment selection for evaluation
VOS and OFS are close but not identical tasks. However, we use VOS benchmarks, since no OFS benchmarks are available. Let us recall that the VOS one is attached to the notion of a primary object of interest moving in the foreground (sometimes, a couple of objects). As a consequence, we have to select the right segments to cope with the binary ground truth of the VOS benchmarks, as usually done for the DAVIS2016, SegTrackV2 and FBMS59 datasets.
Since we deal with multiple-motion segmentation, i.e., $K$ segments, we have to group them into two clusters corresponding to the foreground moving object on one side and the background on the other side. We proceed as in the ST-MS method [4]. However, in our case, we can directly evaluate our LT-MS method in one shot over the predicted segmentation sequence, since there is no temporal linking postprocessing for LT-MS. For evaluations with $K=2$, we do not need to group the masks, so we just set one of the masks as foreground using Hungarian matching over the sequence (alternatively, we could just pick the smallest of the two masks). Regarding the evaluation on the DAVIS2017-motion dataset whose ground truth is not binary, we use the official evaluation script (which perform hungarian matching over the sequence) to associate each predicted segment $k$ with the ground-truth annotations, knowing that we have to consider $K=3$ masks for this experimentation.
A.6 Multi-mask evaluation in FBMS59
We implement the ”Boostrap IoU” score to evaluate multi-mask segmentation in FBMS59, as described in [2] [^3] We use the multi-label ground truth provided by Dong Lao, first author of [2]. The core idea is to match each ground-truth segment to the most likely segment (i.e., with the one of highest IoU). Let us note that this evaluation includes the background segment, since background is not identified in the FBMS59 multi-mask annotation. The algorithm we used for evaluation is detailled in Fig. 6.
Figure 6: Algorithm Evaluation bIoU.
For the linear assignment score, we find the best bipartite sequence-level match between the ground truth and the predictions. Similar to the bIoU evaluation, we compute the score with all the labels of the ground truth. This evaluation is more demanding, because it forces a one-to-one match between the prediction and the masks at the sequence level.
A.7 Repeatability
Figure 7: Model scores depending on initialization. Each dot represents a trained model. The $x$-axis represents the validation loss on the DAVIS 2016 train dataset, and the $y$-axis the performance on the evaluation dataset. The top row includes all models and the bottom row excludes models that diverged during training (validation loss $>0.25$). The bottom row displays the linear relationship between validation loss and evaluation score. The model whose results are reported in the main text is represented as an orange dot.
With the introduction of the transformer decoder, we experimentally found that the convergence of the network depends on the weights initialization, and that the same network and loss configuration can yield different results at test time. In Fig.7, we show that our unsupervised loss on a held-out validation set is a good indicator of the network performance at test time. This is a critical point for model and hyperparameter selection, since we do not have access to the ground truth at training time (with our fully unsupervised scenario), and thus we cannot evaluate model performance. Fig.8 plots the model performance after training as a function of the initialization budget.
In our evaluation experiments, we account for this randomness by training five models with the same set of seeds for all ablations and by reporting the score of the model with the lowest validation loss for each model.
A.8 Training on real-world data
Table 5: Training on synthetic (FT3D) and on real (DAVIS2016) data
Training mode / Dataset DAVIS2016 FBMS59 SegTrackv2
Cut Size 10 120 10 120 10 120
Full Model LT-MS-K4 (FT3D) 74.8 72.4 61.0 58.2 61.3 60.4
Full Model LT-MS-K4 (DAVIS2016) 73.6 72.8 60.0 57.6 62.8 61.3
We have trained our model LT-MS-K4 on the DAVIS2016 training set and evaluated it on DAVIS2016 validation set, FBMS59 and SegTrackV2. Results are collected in Table 5 and compared with those obtained when training our model on the synthetic FT3D dataset. We can observe that performance when training on DAVIS2016 remains globally on par with performance when training on FT3D. Let us note that the volume of data is much smaller when training on DAVIS 2016.
Figure 8: Evolution of the model performance as a function of the initialization budget. Left : the validation loss associated to the best network of each subset. Right : average performance on the test data set associated to this network for each different budget. The filled area correspond to +/- the standard deviation for each curve.
A.9 Influence of the cut size at inference
We have also tested how our LT-MS-K4 method behaves when varying the length of the input flow sequence. Results are plotted in Fig.10 for three datasets, DAVIS2016, FBMS59 and SegTrackV2. As expected, the best performance is obtained for the smallest length (equal to 10) of the input flow sequence. However, performance decreases slowly when the length increases, and remains stable for larger ones. It demonstrates the intrinsic ability of the LT-MS method to achieve accurate and consistent motion segmentation over long periods of the video, which is a unique property. We also did it for LT-MS-K2 as reported in Table 6. We even get slightly better results for DAVIS2016 and SegTrackV2 when the video sequence is processed in one go, i.e., with an infinite cutsize.
Figure 9: Visual assessment of the contribution of the spline-based motion model on the dog video of DAVIS2016 dataset (with $K=4$). From top to bottom: input flows of the sequence; corresponding predicted motion segmentation maps; flows reconstructed with the space-time polynomial motion model; plots over time of the mean-value of the $u$-components of the flows provided per segment by the space-time polynomial motion models; plots over time of the mean-value of the $v$-components of the flows provided per segment by the space-time polynomial motion models; flows reconstructed with the space-time spline-based motion model; plots over time of the mean-value of the $u$-components of the flows provided per segment by the spline-based motion models; plots over time of the mean-value of the $v$-components of the flows provided per segment by the spline-based motion models. Clearly, the space-time polynomial model fails to handle non-uniformly varying motion, whereas the spline-based model does.
Figure 10: Influence of the length of the input flow sequence (also referred to as cut size on the horizontal axis of the plot) on the LT-MS-K4 method for three datasets, DAVIS2016, FBMS59 and SegTrackV2. As expected, the best performance is obtained for the smallest length (cut size equal to 10), but performance decreases slowly when the cut size increases and remains stable for larger ones. On the right, the plot is shown with the full range of values on the ordinate $[0,100]$. On the left, a zoomed-in version of the plot.
Table 6: Results obtained on the three datasets DAVIS2016, FBMS59 and SegTrackV2 with LT-MS-K2 for respectively a cut size of 10 and no cut (the video sequence is processed in one go).
Dataset DAVIS2016 FBMS59 SegTrackV2
Cut Size 10 $\infty$ 10 $\infty$ 10 $\infty$
LT-MS-K2 70.3 70.7 55.3 48.7 58.6 59.3
A.10 Additional visual evaluation
Additional visual results of our LT-MS method are provided in Fig.11. We can observe that the motion segments are globally accurate. Let us recall that the ground-truth is not necessarily available for all the video frames depending on the datasets. Since our method involves several masks, we can properly handle articulated motions (monkey), deal with the presence of several moving objects in the scene (hummingbird), or accomodate motion parallax (libby).
In Fig.12, we collect additional visual results on three videos of the SegTrackV2 dataset, from top to bottom, the birdfall, bird-of-paradise, and bmx videos. Results demontrate the ability of our long-term segmentation method to recover, to some extent, correct segmentation even from flow fields wrongly computed at some time instants.
Figure 12: Illustration of the temporal consistency provided by our LT-MS-K4 method for three examples, the birdfall, bird-of-paradise, and bmx videos of SegTrackV2. For each group, the first row contains the video images with the ground truth overlaid in yellow when available; the second row depicts the corresponding flow fields represented with the HSV code while normalized independently from each other; the third row provides the predicted segmentation.
A.11 Detailed results per videos of the datasets
Hereafter, we report detailed results through tables collecting the evaluation scores obtained by our LT-MS-K4 method for every video of the two datasets DAVIS2016 and Seg-TrackV2 (binary evaluation), and by our LT-MS-K3 model for every video of the DAVIS2017-motion dataset and FBMS59 (multi-mask evaluation). Let us recall that the official evaluation algorithm is not the same for DAVIS2016 and DAVIS2017-motion. The evaluation is done on the whole video for DAVIS2017-motion and is multi-segment, while it is binary and performed frame by frame of the video for DAVIS2016.
A.11.1 DAVIS2016
[…]
,late after line=
,davisnames
[…]
Table 7: Results given for every video of DAVIS2016 dataset. Reported scores per video are the average Jaccard score over frames in the video. The very last row is the average over videos scores. $\mathcal{J}$ is the Jaccard index and $\mathcal{F}$ is the Countour Accuracy. The Mean ($M$) is the average of the scores, the Recall ($O$) is the fraction of frames per video with a score higher than $0.5$, and the Decay ($D$) is the degradation of the score over time in the video.
A.11.2 SegTrackV2
[…]
Rttabular=—l—c—, table head=Video Jacc ($\mathcal{J}$)
, late after line=
, head to column names
[…]
Table 8: Results given for every video of SegTrackV2 dataset. Each reported score is the average Jaccard score over annotated frames in the video. The very last row is the average over all the frames and over all the videos.
A.11.3 DAVIS2017-motion
Table 9: Results given for every video of D17 dataset.
Rttabular=—l—c—c—, table head=Sequence J-Mean F-Mean
, late after line=
, head to column names
[…]
J&FMean JMean JRecall JDecay FMean FRecall FDecay
0.422 0.393 0.387 0.004 0.450 0.437 0.024
Table 10: Results given for every video of DAVIS2017-motion dataset. The very last row is the average score over all the videos for the different criteria.
Table 9: Results given for every video of D17 dataset.

### v3

@@ 5.3 Quantitative and comparative binary evaluation
As shown in Table 2, LT-MS-K2 provides the best results on the two datasets in the category of unsupervised methods based on optical flow only. Besides, we report a repeatability study in the Appendix. OCLR and DivA demonstrate better performance, especially on the SegTrackV2 dataset. However, as aforementioned, OCLR involves additional supervision, while DivA leverages RGB images in its conditional decoder. In addition, DivA, along with MoSeg and CIS methods, takes multi-step flows as input between $t$ and in turn $t+1$, $t+2$, $t-1$, $t-2$, and averages the four corresponding predictions to get the final result.
@@
Failure cases may occur for very complex motion (e.g., ripples on the water), or a particularly strong temporal evolution (as an object approaching the camera from afar quickly and closely). Overall, temporal consistency is properly handled over long periods by our LT-MS method. We want to stress that our method is the only one providing by design a coherent segmentation over the sequence, which is a significant added value for downstream tasks, like tracking or dynamic scene interpretation.
Methods DAVIS2017-motion FBMS59
$\mathcal{J}\&\mathcal{F}$uparrow $\mathcal{J}$uparrow $\mathcal{F}$uparrow bIoU Linear Assign.
@@
Table 3: Multi-segment evaluation, all compared methods using three masks. Evaluation is performed on the video as a whole following the official DAVIS2017 scheme. Reported scores are the average of the individual video scores. The $\mathcal{F}$ metrics focuses on segment boundary accuracy. $\mathcal{J}\&\mathcal{F}$ is the mean of the two. For FBMS59, the evaluation metrics are bootstrap IoU (bIoU) [1] where each ground-truth mask is mapped to the most likely predicted segment (performed at the frame level), and linear assignment that is a bilinear mapping between the ground-truth and the predicted segments at the sequence level (as in the official DAVIS2017 evaluation).
5.4 Multi-segment evaluation
We now focus on the multiple motion segmentation for the DAVIS2017-motion and FBMS59 datasets. Since multiple-motion segmentation is harder than the binary motion segmentation that only involves moving foreground vs background, accuracy scores are expected to decrease for all methods. In Table 3, we report comparative results on the DAVIS2017-motion dataset using the official evaluation procedure, and on FBMS59. As for the other methods, we performed the multiple motion segmentation with three masks ($K=3$), and we name this version LT-MS-K3. To this end, we finetuned the LT-MS-K4 network on the DAVIS2016 training set with now three masks ($K=3$). The resulting performance on DAVIS2017-motion is slightly better than ST-MS and far better than MoSeg. Let us recall that OCLR involves additional supervision. Regarding FBMS59, we report multimask segmentation results with two different metrics. Our LT-MS method outperforms by a large margin DivA [1] that also attempted this multi-segment evaluation.
@@ 5.4 Multi-segment evaluation
Methods DAVIS2016 FBMS59 SegTrackV2
Evaluation basis f s f s f s
LT-MS-K4 75.0 72.4 60.1 58.2 63.4 60.4
ST-MS [2] 71.4 70.2 56.3 55.2 55.3 53.7
EM [29] 77.1 68.0 66.1 51.9 62.5 50.6
Table 4: Accuracy and stability in the evaluation, on DAVIS2016, FBMS59 and SegTrackV2, of our three multiple motion segmentation models, LT-MS, ST-MS and EM, all with $K=4$, on a frame-based (f) basis and on a sequence-based (s) basis for a sequence length of 120. Reported scores correspond to the $\mathcal{J}$ metrics.
5.5 Accuracy vs stability
Since the frame-based quantitative evaluation reported in Section 5.3 cannot fully reflect the contributions of our long-term motion segmentation model, we have carried out complementary experiments. The idea is to somehow cross-validate accuracy and stability. To this end, we compare our three models, the frame-based EM model, the short-term ST-MS model, and the long-term LT-MS model. We proceed in two ways. First, we evaluate all three models on a frame-based basis, then, on a sequence-based basis for a sequence length of 120. By sequence length, we mean the length of the flow sequence taken as input for LT-MS or the output length of the temporal linkage (post-processing stage) for ST-MS.
Results obtained on DAVIS2016, FBMS59 and SegTrackV2 are shown in Table 4 for a binary ground truth. In the Appendix, we explain how we select the segments for the binary evaluation from the set of segments delivered by our multiple motion segmentation methods, when using a number of masks $K>2$. It is not the same procedure as the one used at the end of Section 5.3, which explains the scores obtained for LT-MS-K4 compared to Table 2. Here, we can select the right segments by simply comparing with the ground-truth segments, since it is a question of internal comparison. As expected, the frame-based evaluation favours a frame-based multiple motion segmentation, and overall, the accuracy slightly decreases with the sequence length since it is more demanding to maintain the same labelling, and then to preserve accuracy, over a sequence. In contrast, the sequence-based evaluation highlights the temporal stability brought by ST-MS and with a significant larger margin by LT-MS, in contrast to EM whose performance significantly drops when evaluated on a sequence basis.
@@ 5.6 Qualitative visual evaluation
Visual results are provided for LT-MS-K2 in the Appendix. Fig.4 contains a set of visual results to demonstrate how our LT-MS-K4 method behaves on different situations. We display three result samples obtained on different videos of the benchmarks. We can observe that the motion segments are globally accurate. Since our method involves several masks, we can properly handle articulated motions (people02, bmx), deal with the presence of several moving objects in the scene (people02), separate the rider from the bike or motorbike (bmx, motocross-jump). Since our method enforces temporal consistency, it can deal with errors in optical flow estimation or with objects that momentarily stop moving (bmx).
Additional visual results of our LT-MS-K4 are collected in Fig.5. Again, we can properly cope with articulated motions (monkey), with the presence of several moving objects in the scene (hummingbird, goats01), or accomodate motion parallax (libby). Finally, in Fig.6, we collect visual results on two videos of the SegTrackV2 dataset, from top to bottom, birdfall and bird-of-paradise videos. Results demontrate the ability of our long-term segmentation method to recover correct segmentation even in case of flow fields wrongly estimated for some frames. By the way, this ability was also demonstrated by the greater accuracy (Tables 2 and 4) obtained on SegTrackV2 dataset by LT-MS compared to EM that operates on a frame basis.
Figure 6: Illustration of the temporal consistency provided by our LT-MS-K4 method on two videos of SegTrackV2, birdfall, bird-of-paradise, For each group, the first row contains the video images with the ground truth overlaid in yellow; the second row depicts the corresponding flow fields represented with the HSV code while normalized independently from each other; the third row provides the predicted segmentation.
@@ 6 Conclusion
We have designed an original transformer-based unsupervised method for segmenting multiple motions over a sequence, without any postprocessing. It performs it in one go, which is a more demanding goal, but provides decisive added value for downstream tasks like tracking or dynamic scene interpretation. Our LT-MS method leverages the ELBO framework for the loss function, and fully acknowledges the temporal dimension of the motion segmentation problem. Our method delivers a stable and consistent OF segmentation throughout long video sequences. It introduces at training time B-splines spatio-temporal parametric motion models over space-time volumes, and a loss term expressing temporal consistency over successive masks while taking care of occlusions. Our transformer-based network can be applied at test time to input volumes of any time length. It can accomodate different choices on the number $K$ of masks with a simple fine-tuning step. Experimental results on several datasets demonstrate the efficiency and the accuracy of our LT-MS method by providing competitive results on several datasets. In addition, it is very fast at test time.